Your corporation has the following cash flows: Operating income $250,000 Interest received 10,000 Interest paid 45,000 Dividends received 20,000 Dividends paid 50,000 If the applicable income tax rate is 40 percent (federal and state combined), and if 70 percent of dividends received are exempt from taxes, what is the corporation's tax liability? Select one: a. $ 74,000 b. $ 88,400 c. $ 91,600 d. $100,000 e. $106,500
Answer #1

Taxable Income = Operating Income + Interest Received - Interest Paid + 30% of Dividends Received

(Note: Dividends Paid are Appropriation of Profit. Hence, no adjustment.)

Therefore, Taxable Income = 250000+10000-45000+(0.3*20000) = 221000

Tax Liability = Taxable Income * Tax Rate = 221000*0.4 = $88,400

(If this was helpful then please rate positively. Thank You:)